### v1

# HDN:Hybrid Deep-learning and Non-line-of-sight Reconstruction Framework for Photoacoustic Brain Imaging

Abstract
Photoacoustic imaging (PAI) combines the high contrast of optical imaging with the deep penetration depth of ultrasonic imaging, showing great potential in cerebrovascular disease detection. However, the ultrasonic wave suffers strong attenuation and multi-scattering when it passes through the skull tissue, resulting in the distortion of the collected photoacoustic (PA) signal. In this paper, inspired by the principles of deep learning and non-line-of-sight (NLOS) imaging, we propose an image reconstruction framework named HDN (Hybrid Deep-learning and Non-line-of-sight), which consists of the signal extraction part and difference utilization part. The signal extraction part is used to correct the distorted signal and reconstruct an initial image. The difference utilization part is used to make further use of the signal difference between the distorted signal and corrected signal, reconstructing the residual image between the initial image and the target image. The test results on a PA digital brain simulation dataset show that compared with the traditional delay-and-sum (DAS) method and deep-learning-based method, HDN achieved superior performance in both signal correction and image reconstruction. Specifically for the SSIM index, the HDN reached 0.606 in imaging results, compared to 0.154 for the DAS method and 0.307 for the deep-learning-based method.
{IEEEkeywords}
Deep learning, Photoacoustic imaging, Cerebrovascular disease detection, Non-line-of-sight.
1 Introduction
Cerebrovascular diseases have become one of the most dangerous fatal diseases in the world, and stroke has been listed as one of the top ten causes of death[1]. Timely and effective detection of cerebrovascular stenosis or broken can significantly reduce the mortality of patients.
PAI has emerged as a hybrid imaging modality synergistically combining rich optical contrast with high ultrasonic resolution to provide functional and molecular information from deep tissues. In recent years, PAI has been found to have great potential in human brain imaging[2]. However, the attenuation and scattering by the skull tissue will cause aberration of the PA signal which is a big challenge for brain imaging.
Huang et al. employ detailed subject-specific descriptions of the acoustic properties of the skull to mitigate skull-induced distortions in the reconstructed image. However, this method needs the prior information of CT image data, which are not always available in practice[3]. Liang et al. pointed out that the direct transmission of the shear-wave-converted longitudinal wave is less affected by the skull’s distortion. However, in practice, it is challenging to single out the shear wave signal for image reconstruction [4]. Na et al. proposed layered UBP (L-UBP) method, which can de-aberrate the transcranial PA signal by accommodating the skull heterogeneity into conventional UBP. However, the performance in complex craniocerebral mediators remains to be further explored[5].
In addition, deep-learning-based methods are also popular. Allman et al. proposed the use of convolutional neural networks to identify and remove noise artifacts in photoacoustic signals, which achieved a 100% success rate in classifying both sources and artifacts[6]. Awasthi et al. use deep neural network for super-resolution, denoising as well as bandwidth enhancement of the PA signals collected at the boundary of the domain[7]. Zhang et al. rigorously prepared a photoacoustic digital brain simulation dataset and proposed the use of UNet to correct the PA signal’s distortion. The experimental result shows that the skull’s acoustic aberration can be effectively alleviated after UNet correction, achieving conspicuous improvement in PAT human brain images reconstructed from the corrected PA signal, which can clearly show the cerebral artery distribution inside the human skull[8].
However, the above-mentioned methods focus on how to identify and remove noise or artifacts, few works noticed that within the so-called ”noise” or ”artifacts”, there may be some potentially useful information. Recently, Shen et al. introduced the NLOS imaging into PAI, which treated the skull as an intermediate surface and selected the temporal bone as the imaging window. This method uses reflected PA signal by the inner surface of the skull to expand the imaging field of view and further improve the imaging quality[9].
In this paper, we propose a novel PAI image reconstruction framework, hybridizing deep learning with NLOS imaging. The overview of this paper is arranged as follows. Firstly, we briefly introduced the background in section 2. Then, in section 3, we describe the proposed method in detail. In section 4 and 5, we introduce the simulation detail and present the results of the model, and further discuss the results in section 6. Finally, we draw conclusion in section 7.
2 Background
2.1 Photoacoustic imaging
PAI is based on the PA effect, in which pulsed laser energy is converted into acoustic energy by light absorption. When the laser pulse width meets the thermal confinement and stress confinement, for an ideal point transducer placed at $\vec{r}_{d}$, the detected PA signal can be written as[10]:
[…]
where, $d\Omega$ is the solid-angle element of $\vec{r}$ with respect to the point at $\vec{r}_{d}$, $v_{s}$ is the speed of sound, and $p_{0}(\vec{r})$ is the initial pressure distribution. Then use $x$ and $y$ to represent the PA signal received by the transducers and the initial pressure distribution, respectively. The forward process of PAI in Eq.1 can be simplified as:
$x=Ay$ (2)
where $A$ is a linear operator. The goal of PAI is to solve the above inverse problem by recovering $y$ from the known $x$. DAS is a common and simple analytic reconstruction method. Its basic idea is to backproject the photoacoustic signal at different positions in the same phase[11]:
$y(i,j)=\sum_{n=1}^{N}x(n,t(i,j,n))$ (3)
where $i\in[1,I],j\in[1,J]$ is the size of imaging area, $N$ is the number of the transducers. $t\in[1,T]$ is the delay time of the propagation time of acoustic wave generated at $(i,j)$ to the $n-th$ transducer element.
2.2 Cerebrovascular disease detection
The clinically available techniques for cerebrovascular diseases include magnetic resonance angiography (MRA), computed tomography angiography (CTA), and digital subtraction angiography (DSA)[12]. Though the above methods can reveal clear vascular structure, they also have some drawbacks and limitations. Among them, MRA has problems such as in-plane saturation, intravoxel dephasing, and long acquisition times, which is not advisable for acute stroke patients[13]. Although CTA requires less time, it has problems such as radiation exposure and the risk of contrast nephropathy[14]. DSA is the gold standard for diagnostic cerebrovascular assessment, which is routinely used in the context of neuro-interventional and vascular neurology. Still, it also has some disadvantages, such as higher training demands, complicated pre-procedural preparation, and prohibition to poor renal function patients[15].
Considering that PAI plays an outstanding role in revealing vasculature, it seems that PAI can be competent in the cerebral vascular diagnosis task[16]. In transcranial PAI, the process of receiving signals under ideal conditions is shown in Eq.2. However, due to the existence of the skull, the received signal is multi-scattered and superimposed many times, which will lead to the distortion of the signal:
$x_{dis}(t)=Ay+\sum_{i=1}^{t-1}B_{i}x(i)$ (4)
where $B_{i}$ represents the attenuation and reflection process of the signal $x_{i}$. It is important to note that $B_{i}$ is related to the current time $t$ and does not have time invariance, so it is difficult to decompose $x(t)$ into the target $Ay$ and remainder term by the conventional method.
2.3 Non-line-of-sight imaging
NLOS imaging is a kind of computational imaging technology developed in recent years. As shown in Fig. 1, it collects the light or sound signal $x$ reflected by the intermediate highly-scattering surface through the camera or transducer, to generate the image of the target $y$ beyond the line of sight. At present, NLOS has been widely used in security monitoring, rescue missions, medical imaging, and other fields[17].
Figure 1: Illustration of NLOS PA imaging of vessels in the human brain.
The imaging process of NLOS is similar to the Eq.4, however, due to the obstruction, the signal $Ay$ transmitted by the target cannot be directly received. We receive the signal reflected from the intermediate plane instead:
$x_{ref}(t)=B_{t}Ay+\sum_{i=1}^{t-1}B_{i}x(i)$ (5)
this is similar to the second item in the Eq.4. To make an analogy, if we treat the skull as an intermediate surface and treat the received signal (in Eq.4) as the sum of the target signal (in Eq.2) and the reflected signal (in Eq.5), we hypothesize that a hybrid utilization of both direct PA signal and NLOS signals will help to improve the quality of the reconstructed image.
3 Methods
As shown in Fig. 2, we propose an image reconstruction framework named HDN inspired by NLOS. It consists of the signal extraction part and the difference utilization part. A detailed description of the proposed method will be introduced in the following sections.
Figure 2: The workflow of HDN. The arrows represent the models used, and the symbols in the circle represent subtraction or addition operations.
3.1 Signal extraction part
In transcranial PA imaging, our primary targets are cerebral vessels. However, the acoustic wavelength of the PA signal generated by microvessels is comparable to the size of skull pores, resulting in Mie scattering rather than Rayleigh scattering. This scattering mode will generate a side-lobe signal, and the superimposition of the main lobe and side lobes complicates the PA image reconstruction process[18].
To alleviate PA signal distortion due to the skull tissue, in the signal extraction part, we propose to use the deep-learning-based method to extract a clean signal from the distorted signal and reconstruct the initial image. More specifically, we notice that the collected distorted signal ($x_{dis}$) consists of the target signal ($x_{tar}$) and the reflected signal ($x_{ref}$):
$x_{dis}=x_{tar}+x_{ref}$ (6)
Therefore, we first use the UNet model[19] to extract the corrected signal from the distorted signal:
$x_{cor}=\mathop{\arg\min}\limits_{x_{dis}}\left\|x_{tar}-U_{1}(\theta,x_{dis})\right\|_{2}^{2}$ (7)
where $U(\theta,\cdot)$ denote the UNet map with parameter $\theta$. The UNet structure used in this work is shown in Fig. 3. Then we use the DAS algorithm to reconstruct the initial image. So in the signal extraction section, our goal is to get the initial image from the mixed signal:
$y_{init}=DAS(U_{1}(\theta_{1},x_{dis}))$ (8)
Figure 3: The structure of UNet used in this paper. The numbers in the box represent the kernel size and the number of output channels, respectively.
3.2 Difference utilization part
Previous work only focuses on extracting the corrected signal from the distorted signal. However, it deserves noting that these removed distortions still contain some useful physical information. Inspired by NLOS, we further add the difference utilization part to utilize the removed reflected signal. Specifically, we treat the skull as the reflecting surface, and the difference between the mixed signal and the target signal as the reflected signal.
It is worth noting that the reflected signal experienced multiple reflections and scattering, which makes it difficult to extract features from the signal domain. Therefore, different from the signal extraction part, we first use the DAS algorithm to reconstruct the reflected image:
$y_{ref}=DAS(x_{dis}-x_{cor})$ (9)
then use UNet to learn the residual between the initial image and the target image:
$y_{res}=\mathop{\arg\min}\limits_{y_{ref}}\left\|(y-y_{init})-U_{2}(\theta_{2},y_{ref})\right\|_{2}^{2}$ (10)
By adding the residual image and the initial image, the final imaging result can be obtained:
$\hat{y}=y_{init}+y_{res}$ (11)
4 Simulation detail
4.1 Photoacoustic Digital Brain Dataset
The deep-learning-based approach is a data-driven approach that requires large amounts of data to be trained to get the desired results. However, PAI as a newly developed imaging technology, has not been able to obtain a large amount of clinical data to train neural networks. Therefore, we trained and tested the proposed method using the PA digital brain dataset presented in [8].
Figure 4: The preparation process of PA digital brain dataset[8].
This dataset has a rigorous preparation process. Firstly, T1-weighted 3D MRA images from the IXI dataset from Hammersmith Hospital were used to generate a 2D numerical model of the human brain. By superimposing the layers of the scalp, skull, and other tissues, it can generate a pseudo-3D human brain numerical model size of $12\times 256\times 256$. Then, the optical simulation of the human brain numerical model was carried out by MCXLAB[20], obtaining corresponding optical fluence. The initial pressure distribution is obtained by multiplying the optical flux with the corresponding optical absorption coefficient. Finally, the initial pressure distribution and its removal of skull and scalp tissue were respectively imported into the k-Wave toolbox for acoustic simulation[21], generating distorted signal and target signal. The whole data preparation process is shown in Fig. 4.
4.2 Model Training Details
We trained the HDN using the dataset presented in [8], which contains 180 sets of $256\times 3001$ distorted signal, target signal, and the corresponding $256\times 256$ initial pressure distribution. To maintain dimensional consistency after each pooling operation, we extend the signal to $256\times 3008$ by adding zeros to the end of each channel. We randomly divided the data into 144 training sets, 18 validation sets, and 18 test sets.
In order to better enable the network to learn various features, we carry out data augmentation processing on the training sets. Since the transducers in the acoustic simulation are set to a circular shape and each channel signal comes from the corresponding transducer, we divide the signal evenly into 8 segments, each containing 32 channels. We then select 16 channels from each segment in turn and connect the selected data to get a $128\times 3008$ sampled signal. By overlapping the channels in each segment, we can get 17 different combinations of sampled signals.
We used the Pytorch deep-learning framework to simultaneously train UNet1 and UNet2 on a NVIDIA RTX4090 GPU with 24GB memory. The batch size was 1 due to the data augmentation. The optimizers of the two neural networks were Adam, and the learning rates were 0.001. The loss function is the mean square error (MSE), which is calculated by the following formula:
$\mathrm{MSE=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y_{i}})^{2}}$ (12)
where $y$ represents the truth value and $\hat{y}$ represents the output of the neural network. In the UNet1, we calculate the MSE between $x_{tar}$ and $x_{cor}$, and in the UNet2, we calculate the MSE between $(y-y_{init})$ and $y_{res}$.
5 Results
5.1 Correction and Imaging Results
We did the same data augmentation for the test sets and tested them on the trained model, getting the output and doing the corresponding inverse transformation.
Three samples of normalized target PA sinograms, distorted PA sinograms, and UNet1 corrected PA sinograms are shown in Fig. 5. It shows that the UNet1 effectively separates the clean PA sinograms from the distorted PA sinograms, although there is a certain lack of detail (after $120\mu s$), this does not greatly affect the image quality.
Figure 5: Three samples of normalized PA sinograms. The first column shows the target PA sinograms. The second column is the distorted sinograms used for the inputs of HDN. The UNet1 corrected PA sinograms are in the last column.
The imaging results of the same samples are shown in Fig. 6, compared with the traditional method (DAS) and deep-learning-based method[8] (UNet+DAS, thus the initial image), the HDN has achieved obvious improvement in both image contrast and resolution of cerebrovascular imaging, especially in tiny blood vessels. There is even a slight improvement in the distinction between gray/white matter. Furthermore, the difference between the initial image and the HDN results demonstrated the importance of the difference utilization part, validating the value of multi-reflected PA signals, which are conventionally regarded as ”noise” or ”artifacts”.
Figure 6: Three samples of normalized imaging results. The first column is the initial pressure distribution, which is the imaging target. The second column is the imaging results of the distorted signal directly reconstructed by DAS. The third column is the initial images in the signal extraction part. The last column is the HDN imaging results.
To better quantify the results, the following evaluation indexes were further used to evaluate the performance of the results:
(1). Structural Similarity Index (SSIM), a higher value indicates a better quality for estimated image, which is simply defined as[22]:
$\mathrm{SSIM}(y,\hat{y})=\frac{\left(2\mu_{y}\mu_{\hat{y}}+C_{1}\right)\left(2\sigma_{y\hat{y}}+C_{2}\right)}{\left(\mu_{y}^{2}+\mu_{\hat{y}}^{2}+C_{1}\right)\left(\sigma_{y}^{2}+\sigma_{\hat{y}}^{2}+C_{2}\right)}$ (13)
where $\mu_{y}$ is the local means of $y$, and $\sigma_{y}$ is the standard deviation of $y$. The default values are: C1 equals to 0.01, C2 equals to 0.03.
(2). Peak Signal-to-Noise Ratio (PSNR) is a conventional metric of the image quality in decibels (dB):
$PSNR=20\log_{10}\left(\frac{MAX_{I}^{2}}{\sqrt{MSE}}\right)$ (14)
where $MAX_{I}^{2}$ is the maximum possible pixel value for the image. If each pixel is represented by an 8-bit binary, then the value is 255.
Table 1: Quantitative evaluation results, the first two rows are signal evaluation results, and the last three rows are image evaluation results.
SSIM PSNR MSE
Distorted
0.479 19.913 0.0106
Corrected
0.863 28.524 0.0017
DAS 0.164 11.692 0.0683
UNet1+DAS 0.317 15.261 0.0302
HDN 0.710 22.436 0.0072
We calculated the evaluation indexes between the target signal and the distorted signal before and after signal correction, as well as the evaluation indexes between the target image and the imaging results by different methods. As shown in Tab.1, we found that after correction, the signal has been improved in all the evaluation indexes, especially in the SSIM. Similarly, in terms of imaging results, compared with the traditional analytical method (DAS) and deep-learning-based method (UNet+DAS), the HDN greatly improves imaging quality in three evaluation indexes.
5.2 Ablation Study
Although the difference between the initial image and the HDN results can demonstrate the rationality of the difference utilization part, we still need to explain more. As shown in Fig.7, it is a common step to denoise the PA signal and then reconstruct the initial image in most PA imaging tasks, the same as in the signal extraction part. But inspired by NLOS, we found that there may be some useful information in the discarded noise. In the difference utilization part, we get the reflected signal by subtracting the distorted signal and the corrected signal:
$x_{ref}=x_{dis}-x_{cor}$ (15)
then we prove the above point by using DAS to reconstruct the reflected image. It is worth noting that we do not directly add the initial image to the reflected image. In fact, it is no different from imaging directly using DAS due to the DAS algorithm is a linear transformation:
$y_{das}=DAS(x_{dis})=DAS(x_{cor})+DAS(x_{ref})$ (16)
this will lead to a worse result as shown in the second column in Fig.6. So to make better use of the information in the reflected image, we use UNet2 to learn the residual map between the initial image and the target as a kind of compensation. As shown in Fig.7, the residual image showed a clearer vascular structure compared with the reflected image. Although it seems unreasonable that the residual image is better than the initial image, this is because UNet2’s residual mapping requires the initial image and the target image as prior information, and it doesn’t match the target image numerically.
Figure 7: Three samples of normalized intermediate image in the HDN framework. The first column is the imaging target. The second column is the Initial image in the signal extraction part. The last two columns are the reflected image and residual image respectively in the difference utilization part.
6 Discussion
As shown in Fig. 8, in order to further analyze the distribution of signal correction and image reconstruction results under different conditions, we adopted the kernel density estimation (KDE) method to draw the distribution of the three samples. KDE is a non-parametric estimation method that plots the probability density function of the data, thereby visually showing the distribution of the data over the entire range.
Figure 8: Kernel density estimate (KDE) diagrams of signals and images. (a): The KDE comparison diagram of the three signal samples in Fig. 5. (b): The KDE comparison diagram of the three image samples in Fig. 6.
Fig. 8(a) shows the density distributions of the three signal samples before and after correction. We found that the distribution of the distorted signal in the high-frequency part is much larger than the target, which is caused by signal multi-reflection and superposition. The corrected signal recovers the high-frequency distribution very well, which is the cerebrovascular we are interested in.
Fig. 8(b) shows the density distributions of three PA images under different reconstruction methods. Since the proportion of the imaging regions of interest is small, and most of the values of the regions of no interest are close to or equal to 0, the highest peak can well reflect the imaging of low-frequency details. We found that HDN corrects the position and shape of peaks better than the traditional method and deep-learning-based method. This means that the HDN removes more artifacts and noise so that we can better focus on the parts of interest.
7 Conclusion
In this paper, we propose a PA image reconstruction framework named HDN, which is inspired by NLOS and consists of the signal extraction part and the difference utilization part. The signal extraction part corrects the distorted signal and reconstructs an initial image. The difference utilization part is used to make further use of the signal difference and reconstruct the residual image between the initial image and the target image. HDN is tested on the photoacoustic digital brain simulation data set, and the test results show that compared with the traditional method and deep-learning-based method, HDN has good performance in both signal correction and image reconstruction.